# OmniFusion: Simultaneous Multilingual Multimodal Translations via Modular Fusion

### v1

@@ Appendix A Appendix
A.1 Layer Analysis, Zero-shot ability and APE - AudioTower
Given the strong APE results, we additionally explore whether the model can correct outputs from different ASR systems. Using Whisper (Radford et al., 2023) for ASR, we evaluate the fused APE model on Whisper+Tower outputs and report results in Table˜5. When Whisper+Tower produce the strongest cascaded baseline, using them as hypotheses likewise yields the highest APE performance—highlighting that Speech-Aware APE can be beneficial not only for E2E setups but also for offline cascaded scenarios.
@@ A.2 Prompt Building
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 89.66 2172 132 68 2372
@@
End-to-End (Direct ST)
OmniFusion
Mid Fusion 88.44 2126 157 74 2357
@@
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 86.03 850 932 51 1833
@@
End-to-End (Direct ST)
OmniFusion
Mid Fusion 82.80 754 1065 93 1912
@@
Cascade Systems (ASR → MT)
Omni + SeedX
No FT 81.94 615 1255 104 1974
@@
End-to-End (Direct ST)
OmniFusion
Mid Fusion 79.47 462 1343 50 1855
@@
Table 9: Language codes used in paper following Cheng et al. (2025).

### v2

@@ Appendix A Appendix
Dataset # Examples Language Modality Task
M3AV 122k en Speech, Image ASR, OCR
M3AV - Aug 122k
zh, fr, de, it,
ja, ko, pt, ru,
es, vi, ar, cs,
hr, da, nl, fi,
hu, id, ms, nb,
no, pl, ro, tr
Speech, Image ST, OCR
Europarl 260k*
de, es, fr, it, nl,
pl, pt, ro
Speech ST
Covost2 de, ja, tr, zh Speech ST
Coco-Captions 87k cs, de, fr Image TIT
Table 6: Overview of data sources collected for E2E training. * We sample together 260k from the concatenated data uniformly across the language pairs. All the speech and text data are only available with English as the source language. Language code mapping is presented in Table˜14.
Model Time(s) Image XCOMET-XL
Omni + SeedX FT 2.96 w/o 89.75
OmniFusion 1.98 w/o 88.09
OmniFusion SC 3.78 w/o 88.44
Omni + SeedX FT 3.85 w/ 90.18
OmniFusion 3.15 w/ 89.90
OmniFusion SC 4.82 w/ 90.05
Table 7: Quality vs latency tradeoffs in the offline setting for English $\rightarrow$ German on the MCIF test set. SC indicates Self-Cascade mode.
@@
A.1 Impact of Training Tasks
[…]
W/o OCR 87.12 82.42 79.75
W/o (OCR + Self-Cascade) 89.19 84.35 82.14
All Tasks 89.90 86.79 82.03
All Tasks (Self-Cascade Inference) 90.05 87.20 82.47
Table 9: Ablation study of training tasks on MCIF test set.
To analyze the contribution of the training tasks proposed in Section 2.3, we train two ablation models: one without the OCR task and another without both OCR and self-cascading tasks. We evaluate all models using direct inference on the MCIF test set and report results in Table 9. The results reveal several insights. Removing the OCR task leads to the weakest performance across all language pairs, likely because the model cannot leverage the intermediate ASR step to improve direct translation. Removing both OCR and self-cascading tasks yields moderate performance, as this configuration cannot exploit self-cascading at inference time. In contrast, training on all tasks achieves the best results, particularly when combined with self-cascading inference, demonstrating that our multi-task training strategy effectively leverages both visual information and intermediate ASR outputs without causing negative interference.
@@ A.2 Prompt Building
A.3 Layer Analysis, Zero-shot ability and APE - AudioTower
Given the strong APE results, we additionally explore whether the model can correct outputs from different ASR systems. Using Whisper (Radford et al., 2023) for ASR, we evaluate the fused APE model on Whisper+Tower outputs and report results in Table˜8. When Whisper+Tower produce the strongest cascaded baseline, using them as hypotheses likewise yields the highest APE performance—highlighting that Speech-Aware APE can be beneficial not only for E2E setups but also for offline cascaded scenarios.
@@ A.3 Layer Analysis, Zero-shot ability and APE - AudioTower
Omni FT (Direct ST) LoRA FT 85.90 2047 166 173 2386
Cascade Systems (ASR → MT) Omni + SeedX No FT 89.66 2172 132 68 2372
@@
End-to-End (Direct ST) OmniFusion Mid Fusion 88.44 2126 157 74 2357
@@
Omni FT (Direct ST) LoRA FT 80.88 727 1018 54 1799
Cascade Systems (ASR → MT) Omni + SeedX No FT 86.03 850 932 51 1833
@@
End-to-End (Direct ST) OmniFusion Mid Fusion 82.80 754 1065 93 1912
@@
Omni FT (Direct ST) LoRA FT 80.13 463 1372 44 1879
Cascade Systems (ASR → MT) Omni + SeedX No FT 81.94 615 1255 104 1974
@@
End-to-End (Direct ST) OmniFusion Mid Fusion 79.47 462 1343 50 1855
@@
Table 13: Language codes denoted in our paper following Cheng et al. (2025).
@@
Image SeedX Output OmniFusion Output Reference
[Uncaptioned image] AUSGANG AUSFAHRT AUSFAHRT
[Uncaptioned image] AUSGANG AUSGANG AUSGANG
Table 16: Example illustrating OmniFusion disambiguates source sentence with image context. The word ’EXIT’ when translated to German can be ’AUSFAHRT’ (as vehicle exit) or ’AUSGANG’ (as person exit) and needs the image for disambiguation.